# A 2$\times$2 quantum dot array in silicon with fully tunable pairwise interdot coupling

### v2

@@ Abstract
Recent advances in semiconductor spin qubits have achieved linear arrays exceeding ten qubits. Moving to two-dimensional (2D) qubit arrays is a critical next step to advance towards fault-tolerant implementations, but it poses substantial fabrication challenges, particularly because enabling control of nearest-neighbor entanglement requires the incorporation of interstitial exchange gates between quantum dots in the qubit architecture. In this work, we present a 2D array of silicon metal-oxide-semiconductor (MOS) quantum dots with tunable interdot coupling between all adjacent dots. The device is characterized at 4.2 K, where we demonstrate the formation and isolation of double-dot and triple-dot configurations. We show control of all nearest-neighbor tunnel couplings spanning up to 30 decades per volt through the interstitial exchange gates and use advanced modeling tools to estimate the exchange interactions that could be realized among qubits in this architecture. These results represent a significant step towards the development of 2D MOS quantum processors compatible with foundry manufacturing techniques.
@@ I Introduction
Recent research efforts have already demonstrated significant progress in semiconductor-based spin qubits, showcasing high-fidelity operations Mądzik et al. (2022); Xue et al. (2022); Noiri et al. (2022); Mills et al. (2022); Weinstein et al. (2023); Tanttu et al. (2024), high temperature operations Petit et al. (2018); Yang et al. (2020); Petit et al. (2020); Camenzind et al. (2022); Huang et al. (2024), as well as increasing qubit counts from devices made in both academicPhilips et al. (2022) and foundryGeorge et al. (2024) cleanrooms. One key advantage of a silicon-based platform is the ability to leverage the semiconductor advanced manufacturing capabilities to scale up the number of qubits to millions for a utility-scale quantum computer.
@@
Most of these results have, however, been reported in linear qubit arrays with the current record lying at 12 spin qubits measured in a linear 12-dot device George et al. (2024). Advancement from 1D to 2D arrays of quantum dots is necessary for the development of universal quantum computing architectures compatible with error correction methods such as surface code Fowler et al. (2012). Recent proposals indicate that a 2$\times$×N quantum dot array is enough for a first implementation of error correction Siegel et al. (2024), considering that one of the rows is integrated by physical qubits while the second row enables entanglement via spin-shuttling.
Two-dimensional arrays of semiconductor quantum dots have been implemented in technologies based on heterostructures in GaAs Mukhopadhyay et al. (2018); Mortemousque et al. (2021), Si/SiGe Unseld et al. (2023); Wang et al. (2024a) and Ge/SiGe Wang et al. (2024b); Borsoi et al. (2024); Hendrickx et al. (2021); Zhang et al. (2024). Whilst there have also been demonstrations in silicon MOS, so far they have lacked the tunnel coupling controllability Gilbert et al. (2020); Duan et al. (2020); Ansaloni et al. (2023), which is necessary for optimal Pauli-spin readout and for controlling nearest neighbour entanglement via exchange interactions Tanttu et al. (2024). Modern devices use interstitial exchange gates for this purposeBurkard et al. (2023), however the smaller gate-pitches impose challenges to the fabrication of these devices in academic cleanrooms and in semiconductor foundries Stuyck et al. (2024).
@@
Here, we report the fabrication and measurement of a $\times 2$ quantum dot array in a silicon MOS architecture with integrated interstitial exchange gates in between adjacent quantum dot pairs (Fig. 1(a)). We characterize quantum dot formation at 4.2 K and demonstrate double dot and triple dot configurations in the device. We also demonstrate tunable tunnel coupling between all pairs of nearest neighboring dots. Based on these results and further modeling and simulation, we provide valuable insights into the development of 2D spin qubit architectures.
@@ II Device Design and Fabrication
Figure 1(a) is a scanning electron micrograph (SEM) of a device nominally identical to the one measured in this work. The device consists of a four-layer aluminum gate stack Angus et al. (2007); Lim et al. (2009a), fabricated using electron beam lithography (EBL), thermal deposition of Al and lift-off, on top of thermally-grown high-quality SiO2 on a natSi substrate. Aluminum oxide, AlxOy is formed at each layer to electrically isolate the subsequent layer of Al gates. Figure 1(b) shows the schematic of laying different gates at each layer forming a gate-stack of 2$\times$2 quantum dots with a single-electron transistor (SET). Starting from layer 1 (yellow), we define quantum dot confinement gates (LCB and RCB), SET barrier gates (SLB and SRB), and a barrier gate between dot 2 and dot 3 (J23). Layer 2 (red) consists of plunger gates (P1, P2, P3 and P4), reservoir gate (RG) and an SET top gate (ST). Layer 3 (blue) comprises barrier gates (J12 and J34), reservoir barrier gates (RESB) and an SET barrier gate (SETB). The uppermost layer (purple) consists of a barrier gate (J41) and an on-chip microwave line (ESR). The ESR line is not used in this work. The Al metal thickness on Layers 1, 2, 3 and 4 are 16, 28, 29 and 100 nm, respectively.
In this design, we define all four plunger gates in the same layer but J-gates are defined in different layers. This allows us to investigate the influence of J-gate on interdot tunnel coupling based upon its geometry. Figure 1(c) illustrates the cross-sectional schematics of the possible J-gate geometries. From previous transmission electron microscopy (TEM) analyses of multiple device geometries Lim et al. (2009b); Tanttu et al. (2024); Cifuentes et al. (2024), we see the following trends in gate layer stacking. As shown in Fig. 1(c)(i), J12 (also J34) is fabricated in the $\sim$20 nm gap between the adjacent plunger gates and hence forming a slightly rounded profile on the gate oxide. In comparison, J41 in Fig. 1(c)(ii) is formed with a slightly pointed profile on top of the gate oxide. On the other hand, J23 is fabricated in layer 1, making it a relatively larger barrier gate ($\sim$30 nm). The geometry and position of the J-gates will have a strong influence on the conduction band profiles, as indicated in Fig. 1(c).
@@ III Double Dot Charge Characteristics
In this section, we present the electrical measurement of the 2D array of quantum dots sensed via an SET charge sensor at 4.2 K. We use a gate-pulsed lock-in charge sensing technique Yang et al. (2012) to characterize the charge transitions and electron occupancy of each quantum dot in isolated-mode operation Eenink et al. (2019); Yang et al. (2020). In isolated mode, the RESB and RG gate voltages are reduced to below threshold after the dots are loaded with the desired number of electrons in order to pinch off the tunnel rates to the electron reservoir.
@@
Figure 4: Modelling of exchange coupling in the 2$\times$2 dot array. (a) Electron potential simulation using COMSOL for different double quantum dot pairs as indicated in the inset cartoons and schematic. The electric potentials are simulated with varying J-gate voltage. (b-d) Path integral Monte Carlo simulations of: b) Exchange couplings as a function of J-gate biasing. $V_{J-P}=V_{J}-V_{P}$, where $V_{P}$ is the mean bias of the plunger gates defining the dots. (c) Tunnel control rate by each J-gate in dec/V (d) Exchange control rate by each J-gate in dec/V.
@@ IV Inter-dot tunnel rate measurements
The variations in the J-gate controllability of the tunnel rate in Fig. 3(d) provide an indication of their potential effectiveness in tuning entanglement between spin-qubit pairs formed in the same dots. The exchange rate $J$ is related to the tunnel rate $t$ as
$J\approx\frac{-4t^{2}}{U},$ (1)
where $U$ denotes the Coulomb repulsion. The dependence of both variables under a bias from the exchange gate $V_{\mathrm{J}}$ follows an exponential trend, such that
$\frac{d\log_{10}(J)}{dV_{J}}=2\frac{d\log_{10}(t)}{dV_{J}}.$ (2)
From the tunnel control rates $\frac{d\log_{10}(t)}{dV_{J}}$ measured in Fig. 3(d) of 2.5 dec/V for J41, 10.4 dec/V for J34, and 30.2 dec/V for J23, we can expect exchange rates of 5, 21 and 60 dec/V respectively. Typically an exchange control rate $\frac{d\log_{10}(J)}{dV_{J}}$ higher than $\sim$ 8 dec/V is desirable in order to be able to switch entanglement on and off despite the presence of random sources of variability (e.g. charge traps, Si/SiO2 roughness, etc.)Cifuentes et al. (2024). Under these assumptions, only J41 would fall below that threshold for a suitable J-gate design.
Exchange Gate J23 J12 J34 J14
Metalization Layer 1 3 3 4
Measured Tunnel C. R. 30 10.4 - 2.5
Expected Exchange C. R. 60 21.8 - 5
Simulation Tunnel C. R. 16 4 3 0.5
Simulation Exchange C. R. 31 7 4 1
Table 1: Comparison of measured and simulated control rates on each J-gate. Units are in decades per volt. The experimental control rate is projected from the measured control rate using equation (2).
V Exchange Modelling
The key feature inducing the variations in the tunnel rate control (and the expected exchange rates) is that the J-gates are defined in different metalization layers (See Fig. 1(b)). Each additional metal layer is encapsulated by a thin film of aluminum oxide, which increases their separation from silicon substrate where the dots are formed, thus systematically reducing their effectiveness.
We support this statement with finite element simulations of the electrostatic potential in Comsol Multiphysics, following the methods presented in Ref.[(32)]. We show in Fig. 4(a) that J23 (in layer 1) has the strongest control of the inter-dot potential, as it is the closest to the Si/SiO2 interface, while J14 (in layer 4) is the least effective. We simulate tunnel and exchange rates with a path integral Monte Carlo approach Cifuentes et al. (2023) (see Fig. 4(b--d)), observing a similar trend as the tunnel rates measured in Fig. 3(d). Figures 4(c--d) clearly show that both tunnel and exchange control rates decrease with the metalization layer.
Despite having a good qualitative agreement, the simulated tunnel control rates are significantly smaller than in experiments (Table 1). This is most likely due to limitations in reproducing the device geometry accurately (gate shape, granularity, variations in the oxide thickness, etc.). Modeling could be improved for future studies by carefully analyzing these features with TEM images of the transversal cuts of the device and integrating them into the 3D digital model Cifuentes et al. (2024).
VI Triple Dot Charge Characteristics
@@
VII Conclusion and Outlook
We have fabricated a 2D array of silicon MOS quantum dot and demonstrated pairwise double-dot and triple-dot arrangements in isolated mode at 4.2 K. With the four-layer gate stack, the device can be operated in the single-electron occupancy regime and possesses excellent tunability. The experiments offer important learnings on (i) the design of J-gate and its tunnel-rate controllability; (ii) the position of SET with respect to dots and its sensitivity and (iii) the characterisation of tunnel rates and its relation to exchange rates. The simulation techniques employed can help with future device design optimisation, ensuring high-level exchange controllability for qubit entanglement. These findings highlight the potential of silicon MOS quantum dots as versatile platforms with excellent controllability and connectivity, paving the way for advancements in quantum computing and quantum simulation.

### v3

@@ Abstract
Recent advances in semiconductor spin qubits have enabled linear arrays with more than ten qubits. Scaling to two-dimensional (2D) arrays is essential for fault-tolerant implementations but introduces significant fabrication challenges due to the increased density of gate electrodes. Moreover, implementing two-qubit entanglement control requires the addition of interstitial exchange gates between quantum dots in this dense gate structure. In this work, we present a 2D array of silicon metal-oxide-semiconductor (MOS) quantum dots with tunable interdot coupling between all neighboring dots. Characterized at 4.2 K, the device exhibits exceptional tunability, supports the formation and isolation of both double-dot and triple-dot configurations, and achieves tunnel coupling control spanning up to 30 decades per volt. These results provide critical technical feedback and a foundational benchmark for advancing MOS spin qubit technology into the 2D regime.
silicon, quantum dots, MOS, tunability, tunnel couplings
Keywords: silicon, quantum dots, MOS, tunability, tunnel couplings
@@ I Introduction
Recent research efforts have already demonstrated significant progress in semiconductor-based spin qubits, showcasing high-fidelity operations [1, 2, 3, 4, 5, 6], high temperature operations [7, 8, 9, 10, 11], as well as increasing qubit counts from devices made in both academic[12] and foundry[13] cleanrooms. One key advantage of a silicon-based platform is the ability to leverage the semiconductor advanced manufacturing capabilities to scale up the number of qubits to millions for a utility-scale quantum computer.
@@
Most of these results have, however, been reported in linear qubit arrays with the current record lying at 12 spin qubits measured in a linear 12-dot device [13]. Advancement from 1D to 2D arrays of quantum dots is necessary for the development of universal quantum computing architectures compatible with error correction methods such as surface code [14]. Recent proposals indicate that a 2$\times$×N quantum dot array is enough for a first implementation of error correction [15], considering that one of the rows is integrated by physical qubits while the second row enables entanglement via spin-shuttling.
Two-dimensional arrays of semiconductor quantum dots have been implemented using heterostructures in GaAs [16, 17], Si/SiGe [18, 19], and Ge/SiGe [20, 21, 22, 23]. A key feature in the most advanced implementations is the ability to control tunnel coupling, which is essential for optimal spin readout and enabling entanglement of nearest neighbors via exchange interactions. Modern devices incorporate interstitial exchange gates between quantum dots for this purpose. High-fidelity two-qubit gates were achieved shortly after integrating these gates into the devices [24, 25] and their implementation in 2D arrays have led to the first demonstrations of 2D spin qubit quantum processors [22].
High-fidelity spin qubits have also been demonstrated in MOS quantum dots at the Si/SiO2 interface [11, 6]. These devices are fabricated with materials and processes that are standard in the semiconductor industry [26], which opens up substantial advantages for scalability, including the possibility of fabricating spin qubit arrays at semiconductor foundries [27] and co-integrating them with conventional electronics on the same chip [28, 29].
However, a key limitation of MOS spin qubit technology has been its smaller quantum dot size. The pitch of a state-of-the-art MOS quantum dot device is less than half that of equivalent devices in Si/SiGe or Ge/SiGe [30], demanding very fine lithography with a low error margin. The inclusion of interstitial exchange gates further reduces the already narrow gate pitch by half, pushing the limits of academic cleanrooms and semiconductor foundries [26]. Although there have been demonstrations of 2D quantum dot arrays in silicon MOS [31, 32, 33], none have yet demonstrated tunnel coupling control between all quantum dot pairs.
@@
Here, we report the fabrication and measurement of a $\times 2$ quantum dot array in a silicon MOS architecture with tunable tunnel coupling between all quantum dot pairs (FIG. 1(a)). The device is fabricated with electron beam lithography (EBL) using up to four metalization layers to integrate interstitial exchange gates. We characterize quantum dot formation at 4.2 K and demonstrate double dot and triple dot configurations in the device. We also characterize the tunnel rate control of exchange gates growth in different metalization layers. Based on these results and further modeling and simulation, we provide valuable insights into the development of 2D spin qubit architectures.
@@ II Device Design and Fabrication
Figure 1(a) is a scanning electron micrograph (SEM) of a device nominally identical to the one measured in this work. The device consists of a four-layer aluminum gate stack [34, 35], fabricated using EBL, thermal deposition of Al and lift-off, on top of thermally-grown high-quality SiO2 on a natSi substrate. Aluminum oxide, AlxOy is thermally formed at each layer to electrically isolate the subsequent layer of Al gates. Figure 1(b) shows the schematic of laying different gates at each layer forming a gate-stack of 2$\times$2 quantum dots with a single-electron transistor (SET). Starting from layer 1 (yellow), we define quantum dot confinement gates (LCB and RCB), SET barrier gates (SLB and SRB), and a barrier gate between dot 2 and dot 3 (J23). Layer 2 (red) consists of plunger gates (P1, P2, P3 and P4), reservoir gate (RG) and an SET top gate (ST). Layer 3 (blue) comprises barrier gates (J12 and J34), reservoir barrier gates (RESB) and an SET barrier gate (SETB). The uppermost layer (purple) consists of a barrier gate (J41) and an on-chip microwave line (ESR). The ESR line is not used in this work. The Al metal thickness on Layers 1, 2, 3 and 4 are 16, 28, 29 and 100 nm, respectively.
In this design, we define all four plunger gates in the same layer but J-gates are defined in different layers. This allows us to investigate the influence of J-gate on interdot tunnel coupling based upon its geometry. Figure 1(c) illustrates the cross-sectional schematics of the possible J-gate geometries. From previous transmission electron microscopy (TEM) analyses of multiple device geometries [36, 6, 37], we see the following trends in gate layer stacking. As shown in FIG. 1(c)(i), J12 (also J34) is fabricated in the $\sim$20 nm gap between the adjacent plunger gates and hence forming a slightly rounded profile on the gate oxide. In comparison, J41 in FIG. 1(c)(ii) is formed with a slightly pointed profile on top of the gate oxide. On the other hand, J23 is fabricated in layer 1, making it a relatively larger barrier gate ($\sim$30 nm). The geometry and position of the J-gates will have a strong influence on the conduction band profiles, as indicated in FIG. 1(c).
@@ III Double Dot Charge Characteristics
In this section, we present the electrical measurement of the 2D array of quantum dots sensed via an SET charge sensor at 4.2 K. We use a gate-pulsed lock-in charge sensing technique [38] to characterize the charge transitions and electron occupancy of each quantum dot in isolated-mode operation [39, 8]. In isolated mode, the RESB and RG gate voltages are reduced to below threshold after the dots are loaded with the desired number of electrons in order to pinch off the tunnel rates to the electron reservoir.
@@ IV Inter-dot tunnel rate measurements
This variability in the rate of change of the tunnel coupling is closely linked to the device fabrication process, particularly because the J-gates are deposited in different metallization layers. The simulation tools presented in the Supplementary emulate the EBL fabrication process of the 2$\times$2 array, allowing us to incorporate the impact of these complex features in our modelling (See supplementary FIG. S1). Each additional metal layer is encapsulated by a thin film of aluminium oxide, which increases their separation from the silicon substrate where the dots are formed, thus systematically reducing their effectiveness in tuning the interdot potential barrier (see FIG. S2). The screening of the potential by the gates deposited in previous metal layers can also contribute significantly to this decay. As a result, gates positioned in higher metal layers are expected to exhibit reduced tunability of tunnel couplings and exchange interactions. In the supplementary FIG. S3 we present the results of AB-initio simulations for both parameters using with a path integral approach [40, 37]. These results are in good qualitative agreement with the trends observed in Fig. 3d, further supporting the validity of this explanation.
@@
In this experiment, we were not able to assess spin parameters, such as the exchange interaction, because the measurements were performed at 4.2 K on a silicon substrate without nuclear spin purification—both of which are necessary for high-fidelity spin qubit demonstrations [6]. Nevertheless, we can still estimate the exchange control rate, as it is approximately twice the tunnel control rate $\frac{d\log_{10}(J)}{dV_{J}}\approx 2\frac{d\log_{10}(t)}{dV_{J}}$ (See details in the supplementary). From the tunnel control rates $\frac{d\log_{10}(t)}{dV_{J}}$ measured in FIG. 3(d) of 2.5 dec/V for J41, 10.4 dec/V for J34, and 30.2 dec/V for J23, we can expect exchange rates of 5, 21 and 60 dec/V respectively. This allows us to predict if the device will meet the exchange controllability requirements without the need for dilution refrigeration.
Typically an exchange control rate $\frac{d\log_{10}(J)}{dV_{J}}$ higher than $\sim$ 8 dec/V is desirable in order to be able to switch entanglement on and off despite the presence of random sources of variability (e.g. charge traps, Si/SiO2 roughness, etc.)[37]. Under these assumptions, only J41 would fall below that threshold for a suitable J-gate design. It is worth noting that the dot pitch in P1--P2 and P3--P4 is similar to the ones used in Ref. [6], where high exchange rates were achieved to perform two-qubit gates without accumulating J-dots.
A possible pathway to enhance the tunability of the J41 gate is increasing the pitch between the corresponding plunger gates (P1 and P4). See FIG. S4 in the supplementary material. The gate J41 deposited in the last layer would then fall deeper in the gap between P1 and P4 during the deposition, enabling better control of the interdot potential.
V Triple Dot Charge Characteristics
@@
VI Conclusion and Outlook
We have fabricated a 2D array of silicon MOS quantum dots and demonstrated pairwise double-dot and triple-dot arrangements in isolated mode at 4.2 K. With the four-layer gate stack, the device can be operated in the few-electron occupancy regime and possesses excellent tunability. The experiments offer important learnings on (i) the design of J-gate and its tunnel-rate controllability; (ii) the position of SET with respect to dots and its sensitivity and (iii) the characterisation of tunnel rates and its relation to exchange rates. The simulation techniques presented in the Supplementary can help with future device design optimisation, ensuring high-level exchange controllability for qubit entanglement.
This demonstration underscores the feasibility of 2D MOS spin qubit devices. However, further scaling will require adopting technologies that enable the expansion of the array farther into the two dimensions. This could be achieved by implementing alternative readout methods, such as gate-based readout or single-electron-box sensing [41, 42], which occupy less space than current SETs. The in-plane antenna could also be replaced by an off-chip dielectric resonator [43]. Finally, SLEDGE architectures will enable the vertical fan-out of the gates forming the quantum dot array [44, 45]. In this context, the main contribution of the paper is outlining the necessary benchmarks for gate pitch, tunnel coupling controllability, and sensitivity to interdot transitions to achieve a successful demonstration of a 2D spin qubit array in MOS technology.
SUPPORTING INFORMATION
FIG. S1: Simulation process to generate 3D model used for electrostatic simulations. FIG. S2: Electrostatic simulations of the device. FIG. S3: Path integral simulations of tunnel and exchange couplings. FIG. S4: Exchange control rate vs gate pitch.
@@ ACKNOWLEDGEMENTS
Appendix A Device Modeling
The device modeling and simulation consists of three stages:
  1. 1.
    Simulation of the device fabrication process (FIG.5)
  2. 2.
    Finite-element simulation of the electrostatic potential in COMSOL Multiphysics (FIG.6)
  3. 3.
    Ab initio path integral simulations of tunnel couplings and exchange interactions [40] (FIG. 7)
Using this protocol, it is possible to predict the potential impact of changes in the device parameters, such as the gate-to-gate pitch (FIG. 8).
Figure 5: a, Simulation of the device fabrication process described in FIG. 1 of the main text, emulating the thermal expansion of the four metal layers following deposition and subsequent oxide growth. b, Final 3D model imported into COMSOL Multiphysics for electrostatic simulations. The simulation parameters used to generate the 3D geometry are carefully tuned to closely match the physical device observed in SEM images [37] (see FIG. 1a of the main text).
Figure 6: a, Cross section of finite element electrostatic simulations in COMSOL Multiphysics of the potential of the computer model in FIG 5b for each double quantum dot pair (see inset schematics). The electric potentials are simulated with varying J-gate voltage showing distinct levels of effectiveness for each gate.
Figure 7: a-b, Path integral Monte Carlo simulations of a) tunnel couplings and b) exchange couplings as a function of J-gate biasing in the model of FIG. 5b (Example electrostatic potentials in FIG. 6). $V_{J-P}=V_{J}-V_{P}$, where $V_{P}$ is the mean bias of the plunger gates defining the dots (See Ref. [40, 37] for details on the path integral simulation methods and bechmarking with experiments [37] and FCI simulations [40]). c-d, Electrostatic control rate of c) the tunnel coupling $\left(\frac{d\log_{10}t}{dV_{J}}\right)$ and d) the exchange coupling $\left(\frac{d\log_{10}J}{dV_{J}}\right)$ by each J-gate in dec/V. The control rates are in qualitative agreement with the measurements in FIG. 3d of the main text with gate J23 being the most effective and J14 the less effective. We note, however, that the values obtained for the simulated tunnel control rates are significantly lower than experimental results limiting any quantitative comparison between experiments and simulations. This could occur because of the difficulty on reproducing the device geometry accurately (gate shape, granularity, variations in the oxide thickness, etc.) in the absence of TEM images of the transversal cuts of the device as in Ref. [37]. These features are critical to obtain a good estimate of the exchange control.
Appendix B Exchange Control Rate vs Tunnel Control Rate
The exchange control rate can be approximated from the tunnel control rate as
$\frac{d\log_{10}(J)}{dV_{J}}\approx 2\frac{d\log_{10}(t)}{dV_{J}}$ (1)
This equation can be initially understood as the derivative of $J\approx\frac{4t^{2}}{U}$ with respect to $V_{J}$, derived from the Hund-Mulliken model [30]. Although this model may break down under strong interaction regimes, our previous experimental work [6] - where tunnel and exchange rates were measured in the same device - suggests that Equation (1) remains valid in practice. In addition, in FIG. 7c-d, we present ab-initio path integral simulations of tunnel and exchange couplings for the four-dot system [40, 37], which also agree with equation (1). This method has been benchmarked against Full CI simulations  [40] and experimental data [37], and is expected to remain valid at the strongly interacting regime, making it more reliable than the Hund-Mulliken model. This suggests that measuring tunnel-rate controllability at 4 K can serve as an effective predictor of whether a device will meet the exchange controllability requirements across all qubit pairs - without the need for dilution refrigeration.
Figure 8: Simulation of the dependence of the exchange control rate on the inter-dot pitch. This simulation uses a simplified two-dot device model, distinct from the four-dot device shown in FIG. 5b. Although the device architecture differs, the fabrication process is similar, so the observed trend is expected to hold for the four-dot device as well.